OpenStudy (anonymous):

A car accelerates uniformly in a straight line from rest at the rate of 2.8 m/s2 . What is the speed of the car after it has traveled 50 m?

OpenStudy (anonymous):

this is my way of doing this: (v final)^2 = (v initial)^2 + 2as v^2 = 0 + 2(2.8m/s2)50m v^2 = 280m2/2 v = 16.73 m/s

OpenStudy (anonymous):

thank you so much!! :)

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

if you want, can you help me again. Now with the same question above, i need to find how many seconds it took the car to travel in 50m. Im just not getting what equation to use, do i find acceleration first?

OpenStudy (anonymous):

no, you already have acceleration, which is 2.8m/s2. This is my way of doing this. Use the formula: (v final) = (v initial) + at 16.73m/s = 0 + (2.8m/s2)(t) => t = 5.975s

OpenStudy (anonymous):

Thank you! Youre a life saver